# Alignment via LLM-based Embedder

### v2

@@ 4.2 Main Results
Method Data
@@
ProCLIP 30M 88.9 94.1 77.7 72.5 61.1 25.2 62.8 81.5 92.9 57.2 67.8 71.0 (+13.5)
Model Architecture: EVA02-CLIP ViT-L/14
EVA02-CLIP 2B 92.9 98.8 89.8 73.8 88.8 35.1 60.6 93.7 95.1 76.3 78.2 80.3
LLM2CLIP 3M 64.1 96.5 82.6 68.0 29.2 9.0 59.4 48.5 89.8 28.6 56.4 57.5
ProCLIP 3M 82.7 97.9 88.4 73.6 57.6 16.5 63.5 67.6 93.8 45.4 66.8 68.5 (+11.0)
Multilingual Cross-Modal Retrieval. Benefiting from the LLM-based embedder, ProCLIP facilitates multilingual capabilities. As illustrated in Fig. 3, we compare the cross-lingual retrieval performance between LLM2CLIP and ProCLIP on the XM3600 benchmark (Thapliyal et al., 2022). Experiment results demonstrate that our approach achieves superior multilingual performance. This enhancement is attributed to the improved alignment between the CLIP image encoder and the LLM-based embedder.
Zero-Shot Classification. In Tab. 2, we present the zero-shot classification performance on 11 downstream tasks across different data and model scales. We observe that LLM2CLIP significantly compromises the original generalization ability of CLIP. Even when utilizing 30M data points, compared to CLIP’s ViT-B/32, ViT-B/16, and ViT-L/14, the average performance declines by 16.2%, 15.3%, and 18.2%, respectively. Compared to LLM2CLIP, our proposed ProCLIP method achieves significant performance improvements across all experimental conditions. Particularly, with a dataset of 30M samples, ProCLIP enhances the average performance by approximately 10%-13.5%. This notable improvement is primarily attributed to two factors: 1) The representation inheritance process allows the LLM embedder to inherit some knowledge from the original CLIP text encoder. 2) During the contrastive tuning phase, the introduction of a distillation loss as regularization helps to mitigate the forgetting of knowledge throughout the learning process.
Table 3: Robustness performance. The best results are marked in bold.
Method Data Robustness
@@
ProCLIP 30M 63.4 68.0 34.1 86.8 55.7
Model Architecture: EVA02-CLIP ViT-L/14
EVA02-CLIP 2B 72.6 76.4 29.6 92.7 67.9
LLM2CLIP 3M 51.8 50.4 28.8 79.1 50.6
ProCLIP 3M 62.0 66.5 29.3 89.4 59.9
Robustness. To evaluate the robustness of ProCLIP  we report its performance across varying data sizes and model scales in Tab. 3. ProCLIP consistently achieves average improvements of 5.9%-9.3%. Notably, on challenging out-of-distribution datasets like ImageNet-A and ImageNet-R, ProCLIP outperforms LLM2CLIP by over 10 percentage points, highlighting its enhanced ability to handle distribution shifts and complex perturbations. These results demonstrate that ProCLIP not only improves retrieval and classification performance but also delivers robust and reliable results across diverse robustness scenarios, indicating substantial progress in generalization and resilience.
Figure 4: MMVP performance comparison. ProCLIP presents excellent performance.
Method ViT Init LLM Embedder Data ImageNet COCO Flickr30k
@@
ProCLIP CLIP L/14 NV-Embedv2 3M 61.4 64.8 51.7 91.9 81.4
ProCLIP CLIP L/14 Llama3-$C$ 3M 62.5 66.4 51.9 92.8 81.1
Fine-Grained Understanding. Fig. 5 presents the fine-grained vision-language understanding performance on the MMVP benchmark (Tong et al., 2024) using CLIP ViT-L/14. LLM2CLIP improves over CLIP by 2.9%, 5.9%, and 4.4% at 3M, 15M, and 30M data scales, respectively. Our ProCLIP model further advances these results, achieving gains of 3.0%, 2.2%, and 10.4% on the corresponding data scales. These improvements demonstrate that the LLM-based embedder enhances fine-grained semantic discrimination, and the consistent superiority of our method underscores the effectiveness of the progressive alignment strategy.
Comparison with Other Methods. To further prove the effectiveness of ProCLIP, we provide a comprehensive comparison of all recent LLM embedder-based CLIP models, including FLAME (Cao et al., 2025), ShareLock (Ruthardt et al., 2024), LIFT (Yang et al., 2025), SAIL (Zhang et al., 2025a), LiT (Zhai et al., 2022), and our baseline LLM2CLIP (Huang et al., 2024). As shown in Tab. 5, under the same or lower training costs, ProCLIP consistently achieves superior performance across various model sizes. Benefiting from representation inheritance and self-distillation regularization, ProCLIP not only achieves significant performance improvements in In1k classification but also enhances general retrieval capabilities on COCO and Flickr30k.
4.3 Ablation Study
@@
Figure 7: Ablation on different components.
Method Stage 1 Stage 2 IN-1k I2T Avg T2I Avg
@@
62.5 83.0 80.5
Ablation of Different LLM-based Embedder. As shown in Fig. 7, we compare different LLM embedders, including Qwen3-Embedding (8B), GME (7B), NV-Embedv2 (7B) , and Llama3-CC (8B) based on ViT-L/14 with 15M data. Llama-CC achieves the strongest overall performance in both ImageNet zero-shot classification and retrieval tasks. Notably, while different embedders show only minor variations in retrieval performance, they exhibit substantial differences in ImageNet classification accuracy. This suggests that the alignment discrepancy between each LLM embedder’s feature space and the original CLIP space varies considerably, resulting in different degrees of degradation in general capabilities after image-text alignment.
Ablation of Different Components. To further validate the effectiveness of the methods proposed in this paper, we conduct a comprehensive ablation study on various components, as detailed in Tab. 7. Applying instance semantic distillation achieves 58.9% zero-shot accuracy on ImageNet-1k using only text data, indicating successful transfer of CLIP’s textual representation capability to the MLP head. Incorporating the structural alignment loss further improves both classification and retrieval performance by enabling the LLM embedder to capture the global structural geometry of CLIP’s text representation space, beyond point-wise semantic correspondences. After that, image-text contrastive learning significantly boosts retrieval performance but reduces ImageNet-1k accuracy due to image encoder overfitting. Introducing self-distillation mitigates this issue, improving classification accuracy from 59.2% to 62.5% while slightly reducing retrieval gains. Finally, applying structured self-distillation enhances both tasks by stabilizing the image representation space during fine-tuning, preventing excessive overfitting while preserving pretrained knowledge.
@@ 5 Conclusion
In this paper, we propose ProCLIP, a simple yet effective progressive vision-language alignment framework designed to improve the alignment when integrating the CLIP image encoder with an LLM-based embedder. The framework employs a curriculum learning–inspired progressive training strategy: it first aligns the LLM-based embedder’s representation space with the original CLIP text encoder through knowledge distillation, effectively transferring pretrained semantic knowledge. Subsequently, it performs cross-modal alignment between the CLIP image encoder and the LLM-based embedder using image-text contrastive learning regularized by self-distillation to prevent overfitting and preserve pretrained knowledge. To ensure feature-space consistency, a complementary distillation strategy—comprising instance semantic and embedding structure alignment losses—is applied during text distillation and image self-distillation, respectively. Comprehensive experiments across varying data scales and model architectures demonstrate the effectiveness and generality of ProCLIP. We hope that our work offers valuable insights for advancing vision-language alignment. We will release all model weights and code to ensure full reproducibility.
@@ A.1 Training Details
Table 6: Zero-shot image-text retrieval evaluation settings.
@@
Multilingual Retrieval.
We evaluated the multilingual capabilities of our model on XM3600 (Thapliyal et al., 2022). XM3600 contains 3,600 images covering a total of 36 languages, including Arabic (ar), Bengali(bn), Chinese-Simplified (zh), Croatian (hr), Czech (cs), Danish (da), Dutch (nl), English (en),Farsi (fa), Filipino (fil), Finnish (fi), French (fr), German (de), Greek (el), Hebrew (he), Hindi (hi), Hungarian (hu), Indonesian (id), Italian (it), Japanese (ja), Korean (ko),Maori(mi), Norwegian (no), Persian (fa), Polish (pl), Portuguese (pt), Romanian (ro), Russian (ru), Spanish (es), Swedish (sv), Swahili(sw), Thai (th), Turkish (tr), Telugu (te), Ukrainian (uk), and Vietnamese (vi).
Fine-Grained Understanding. We evaluated the fine-grained understanding capability of the VLM on MMVP-VLM (Tong et al., 2024). MMVP-VLM consists of 150 samples in total, testing 9 patterns:
@@
  • L State and Condition: Questions that pertain to the state or condition of an object, such as whether a flag is blowing in the wind or if the ground is wet.
  • Æ Quantity and Count: Questions about the number of objects or features present in the image.
  • , Positional and Relational Context: This aspect refers to the model’s ability to understand the position and relationship of objects or elements within an image in relation to each other and their surroundings.
  • h Color and Appearance: Questions regarding the color of certain objects or elements.
@@
  • k Text: Questions related to text or symbols present in the image.
@@
MLLM benchmarks. We further integrate the fine-tuned vision encoder into LLaVA and evaluate its performance on several MLLM downstream benchmarks, including SEED-Bench (Li et al., 2023a), GQA (Hudson & Manning, 2019), VizWiz (Gurari et al., 2018), PoPE (Li et al., 2023b), TextVQA (Singh et al., 2019), MMBench (Liu et al., 2024b), and VQAv2 (Goyal et al., 2017).
A.3 More results.
Liner Probe. We conduct linear probe evaluations of the model on 11 datasets. As shown in 9, our method consistently achieves superior performance. This advantage stems from our progressive alignment framework, which stabilizes training through two-stage regularization that prevents overfitting in the vision encoder while preserving generalization capability.
Table 7: MLLM(7B) performance under $224^{2}$ image resolution.
@@
ProCLIP 66.4 62.3 44.7 85.9 54.5 65.0 78.1
@@
Method Data Alignment Layer IN-1k I2T Avg. T2I Avg.
CLIP 400M - 74.5 66.9 55.5
ProCLIP 1M $4\timeslinear$ 61.8 81.9 79.7
LLM2CLIP 3M $4\timeslinear$ 52.8 82.0 79.1
ProCLIP 3M $4\timeslinear$ 62.5 83.0 80.5
LLM2CLIP 15M $4\timeslayers$ 54.0 83.2 80.1
ProCLIP 15M $4\timeslayers$ 66.0 84.3 81.2
LLM2CLIP 30M $4\timeslinear$ 56.4 85.8 82.1
ProCLIP 30M $4\timeslinear$ 67.8 86.2 82.6
ProCLIP 30M $12\timeslinear$ 71.5 86.8 82.8
Table 9: Linear Probe performance on 11 datasets.
@@
ProCLIP 3M 95.3 99.5 94.0 81.0 93.9 65.7 85.9 95.4 97.8 99.3 84.5 90.2
@@
A.4 Limitations and Future Works.
Training Efficiency. In our proposed progressive alignment framework, the additional computational overhead in the first stage is relatively small. However, in the second stage, the vision encoder needs to be unfrozen for training, and self-distillation is performed online, which increases the computational cost. In our experiments, the training speed is approximately 0.74× that of the baseline. We consider the following directions to potentially reduce computational overhead:
@@
Fine-grained Visual Alignment. Our proposed ProCLIP is still based on contrastive learning over global semantics. However, aligning local vision patches with textual semantics can enhance the local perception capabilities of the vision encoder, benefiting tasks such as open-vocabulary segmentation and open-vocabulary detection. Further discussions on this topic can be found in works like SigLIP2 (Tschannen et al., 2025) and FG-CLIP (Xie et al., 2025).
More Model Architecture. Our approach replaces the original CLIP text encoder with an LLM-based embedder to enhance multiple capabilities. From another perspective, can we similarly replace the vision encoder in the dual-tower architecture to address limitations in visual representation? For example, CLIP’s image encoder is known to lack locality—could this limitation be mitigated by substituting the image encoder? We plan to explore this direction further in future work.

### v3

@@ 3.1 Preliminary
$\mathcal{L}_{reg}=\sum_{i=1}^{\mathcal{B}}\mathcal{L}_{ins}(v_{i},v_{i}^{*})+\sum_{i<j}^{\mathcal{B}}\mathcal{L}_{struct}(v_{i},v_{j}^{*})$ (5)
where $v_{i}^{*}=\mathcal{E}_{I}^{*}(I_{i}),v_{j}^{*}=\mathcal{E}_{I}^{*}(I_{j})$ denote the vision embeddings derived from the EMA-updated image encoder. The EMA update procedure is defined as follows:
$\mathcal{E}_{I}^{*}=\alpha\mathcal{E}_{I}^{*}+(1-\alpha)\mathcal{E}_{I},$ (6)
where $\alpha$ controls the update rate of the teacher model parameters. The overall loss function of the contrastive tuning stage is defined as $\mathcal{L}_{\text{tune}}=\mathcal{L}_{\text{info}}+\lambda_{reg}\mathcal{L}_{\text{reg}}$, where $\lambda_{reg}$ is a loss weight.
4 Experiments
Method #Data
@@ 4 Experiments
[…]
Method #Data Robustness
@@
[…]
Table 3: Robustness performance. The best results are marked in bold.
4.1 Experimental Setup
Datasets and Benchmarks. For the alignment dataset, we use CC3M (Changpinyo et al., 2021), CC12M (Changpinyo et al., 2021), and YFCC15M (Thomee et al., 2016),combined the high-quality captions from DreamLIP (Zheng et al., 2024). We conduct experiments with data scales of 3M (CC3M), 15M (CC3M + CC12M), and 30M (CC3M + CC12M + YFCC15M) to explore the effects of data scaling. For the benchmark, we perform zero-shot classification on 11 different classification datasets, robustness evaluations on 5 ImageNet variants, retrieval evaluations on 6 datasets, multilingual cross-modal retrieval evaluation on XM3600 (Thapliyal et al., 2022), and fine-grained understanding evaluation on MMVP-VLM (Tong et al., 2024) and SugarCrepe (Hsieh et al., 2023). Regarding the model, we mainly employ three OpenAI pre-trained CLIP models, ViT-B/32, ViT-B/16, and ViT-L/14, to investigate the effects of model scaling. Additionally, we conduct experiments with pretrained EVA02-CLIP (Fang et al., 2023) ViT-L/14 and SigLIP2 Tschannen et al. (2025) SO/14@224 to assess the impacts of different model architectures. For the LLM-based embedder, we primarily use LLaMA3-8B-CC consistent with LLM2CLIP Huang et al. (2024).
Figure 4: Compositional performance on SugarCrepe.
Implementation Details. For the representation inheritance phase, we train for four epochs, followed by another four epochs for contrastive tuning. During training, we employ AdamW (Loshchilov, 2019) as the optimizer, with a learning rate of $1\times 10^{-5}$ and a weight decay of 0.2. The parameters $\beta_{1}$ and $\beta_{2}$ are set to 0.9 and 0.98, respectively. In the first stage, the training batch size is set to 1024, while in the second stage it is increased to 4096. The loss weight $\lambda_{r}eg$ is set at 0.0004. Other training details can be found in the appendix.
Figure 5: MMVP performance.
Method ViT Init LLM Embedder #Data IN-1k COCO Flickr30k
@@ 4.1 Experimental Setup
[…]
4.2 Main Results
Cross-Modal Retrieval. Tab. 1 shows that ProCLIP consistently outperforms LLM2CLIP across all datasets and model scales. In short-text scenarios (e.g., Flickr30k and COCO), ProCLIP achieves substantial gains; notably, with ViT-L/14 (30M), it reaches 95.0% I2T R@1 on Flickr30k, outperforming LLM2CLIP by  2%. This advantage extends to long-text benchmarks (DOCCI, DCI, Urban-1k), where ProCLIP (ViT-B/16, 30M) attains 73.6% I2T and 76.9% T2I R@1 on DCI. Across all training scales (3M to 30M), ProCLIP yields stable improvements, particularly in T2I retrieval. These results validate ProCLIP’s versatility in handling both concise and complex textual scenarios.
Multilingual cross-modal Retrieval. Leveraging the LLM-based embedder, ProCLIP exhibits robust multilingual capabilities. As illustrated in Fig. 3, ProCLIP consistently outperforms LLM2CLIP on the XM3600 benchmark (Thapliyal et al., 2022) across various languages. This superiority is primarily attributed to our progressive alignment strategy, which more effectively bridges the semantic gap between the CLIP image encoder and the multilingual LLM space while preserving pretrained visual priors.
Zero-Shot Classification. Tab. 2 summarizes zero-shot classification performance across 11 downstream tasks. We observe that LLM2CLIP severely compromises CLIP’s inherent generalization, with average performance dropping by 16.2% (ViT-B/32), 15.3% (ViT-B/16), and 18.2% (ViT-L/14) even when trained on 30M samples. In contrast, ProCLIPconsistently achieves substantial improvements over LLM2CLIP across all configurations, yielding a 10.0%–13.5% average gain on the 30M dataset. This superiority stems from the well-crafted alignment curriculum of ProCLIP, which leverages pretrained knowledge to achieve more effective and robust cross-modal alignment.
Robustness. Tab. 3 evaluates the robustness of ProCLIP across various data scales and model architectures, showing consistent average gains of 5.9%–9.3%. Notably, on challenging out-of-distribution (OOD) benchmarks such as ImageNet-A and ImageNet-R, ProCLIP outperforms LLM2CLIP by over 10 percentage points. This substantial margin highlights ProCLIP’s superior capacity to handle distribution shifts and complex perturbations.
@@ 4.2 Main Results
Method Stage 1 Stage 2 IN-1k I2T Avg. T2I Avg.
@@
[…]
Figure 8: Ablation on different components.
Fine-Grained Understanding. Fig. 4 demonstrates the significant potential of ProCLIP in compositional understanding tasks, surpassing the CLIP baseline by an average margin of 11.3%–17.7%. These improvements demonstrate that the LLM-based embedder enhances fine-grained semantic discrimination, and the consistent superiority of our method underscores the effectiveness of the progressive alignment strategy. Fig. 6 presents the fine-grained vision-language understanding performance on the MMVP benchmark (Tong et al., 2024) using CLIP ViT-L/14. LLM2CLIP improves over CLIP by 2.9%, 5.9%, and 4.4% at 3M, 15M, and 30M data scales, respectively. Our ProCLIP model further advances these results, achieving gains of 3.0%, 2.2%, and 10.4% on the corresponding data scales.
Comparison with Other Methods. To further prove the effectiveness of ProCLIP, we provide a comprehensive comparison of all recent LLM embedder-based CLIP models, including FLAME, ShareLock, LIFT, SAIL, LiT, and our baseline LLM2CLIP. As shown in Tab. 6, under the same or lower training costs, ProCLIP consistently achieves superior performance across various model sizes. Benefiting from representation inheritance and self-distillation regularization, ProCLIP not only achieves significant performance improvements in In-1k classification but also enhances general retrieval capabilities on COCO and Flickr30k.
4.3 Ablation Study
Ablation of Different LLM-based Embedder. Fig. 8 compares several LLM embedders (Qwen3, GME, NV-Embedv2, and Llama3-CC) paired with ViT-L/14. Among them, Llama3-CC achieves the best overall results in classification and retrieval. Notably, while retrieval performance remains relatively consistent across embedders, ImageNet classification accuracy varies significantly. This suggests that the alignment discrepancy between each LLM’s feature space and the original CLIP space differs, leading to varying degrees of generalization degradation after training.
Base Model Method IN-1k I2T Avg. T2I Avg.
CLIP ViT-L LLM2CLIP 52.8 82.0 79.1
[…]
EVA02-CLIP ViT-L LLM2CLIP 56.4 83.6 80.9
[…]
SigLIP2 SO/14@224 LLM2CLIP 63.8 84.6 81.9
[…]
Table 4: Ablation study on different base model architectures. All methods use 3M training samples.
Ablation of Different Base Model. Tab. 4 shows ProCLIP excels across architectures, surpassing LLM2CLIP in classification and retrieval. The remaining classification gap relative to vanilla CLIP arises from the latter’s overwhelming data scale, a dominant factor in zero-shot tasks.
Ablation of $\lambda_{reg}$. Tab. 5 presents the sensitivity analysis of the regularization weight for $\lambda_{\text{reg}}$. The results indicate that 4e-4 offers an optimal balance and is thus selected as our final configuration. Furthermore, the empirical evidence suggests that excessively strong regularization hampers alignment, leading to suboptimal results.
Ablation of Different Components. Tab. 8 details the ablation study of our components. First, instance semantic distillation achieves 58.9% ImageNet-1k zero-shot accuracy using text only, confirming the successful transfer of CLIP’s textual priors to the adapter. Adding structural alignment loss further boosts both classification and retrieval by enabling the LLM to capture CLIP’s global geometric structure beyond point-wise semantics. Subsequent image-text contrastive tuning significantly enhances retrieval but compromises classification due to image encoder overfitting. This is mitigated by self-distillation, which recovers classification accuracy from 59.2% to 62.5%. Ultimately, structured self-distillation yields the best results by stabilizing the image representation space and preserving pretrained knowledge, effectively balancing fine-tuning gains with generalization.
$\lambda_{reg}$ IN-1k I2T Avg. T2I Avg.
0 59.2 82.9 80.2
2e-4 60.3 82.8 80.3
[…]
6e-4 62.6 82.6 79.8
1e-3 62.5 81.8 79.5
Table 5: Ablation of $\mathcal{L}_{reg}$
@@ 5 Conclusion
This paper presents ProCLIP, a progressive vision-language alignment framework for integrating CLIP image encoders with LLM-based embedders. Inspired by curriculum learning, ProCLIP employs a two-stage strategy: first, it anchors the LLM-based embedder to CLIP’s textual space via knowledge distillation to inherit pretrained semantic priors; second, it performs cross-modal contrastive tuning while utilizing self-distillation to prevent overfitting. To maintain feature-space consistency, we introduce a complementary distillation scheme incorporating instance semantic and structural alignment losses. Extensive evaluations across various scales and architectures validate the efficacy and generality of our approach.
6 Limitation
[…]
@@ A.1 Training Details
Table 8: Zero-shot image-text retrieval evaluation settings.
Multilingual Retrieval. We evaluated the multilingual capabilities of our model on XM3600 (Thapliyal et al., 2022). XM3600 contains 3,600 images covering a total of 36 languages, including Arabic (ar), Bengali(bn), Chinese-Simplified (zh), Croatian (hr), Czech (cs), Danish (da), Dutch (nl), English (en),Farsi (fa), Filipino (fil), Finnish (fi), French (fr), German (de), Greek (el), Hebrew (he), Hindi (hi), Hungarian (hu), Indonesian (id), Italian (it), Japanese (ja), Korean (ko),Maori(mi), Norwegian (no), Persian (fa), Polish (pl), Portuguese (pt), Romanian (ro), Russian (ru), Spanish (es), Swedish (sv), Swahili(sw), Thai (th), Turkish (tr), Telugu (te), Ukrainian (uk), and Vietnamese (vi).
Fine-Grained Understanding. We evaluated the fine-grained understanding capability of the VLM on MMVP-VLM (Tong et al., 2024) and SugarCrepe. MMVP-VLM consists of 150 samples in total, testing 9 patterns:
REPLACE SWAP ADD
Object Attribute Relation Object Attribute Object Attribute
# negative captions 1652 788 1406 246 666 2062 692
Table 9: Number of hard negative captions of all types in SugarCrepe.
Method Flickr30k COCO ShareGPT4V Urban-1k DOCCI DCI Avg.
I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I I2T T2I
CLIP 86.6 64.6 57.2 36.4 78.0 68.7 68.4 56.0 65.8 63.1 45.4 43.9 66.9 55.5
Long-CLIP 89.7 75.7 62.5 46.4 93.3 92.1 82.4 86.2 66.4 78.6 57.2 64.0 75.2 73.8
[…]
Table 10: Compared to Long-CLIP on multiple cross-modal retrieval datasets.
@@
  • State and Condition: Questions that pertain to the state or condition of an object, such as whether a flag is blowing in the wind or if the ground is wet.
  • Quantity and Count: Questions about the number of objects or features present in the image.
  • Positional and Relational Context: This aspect refers to the model’s ability to understand the position and relationship of objects or elements within an image in relation to each other and their surroundings.
  • Color and Appearance: Questions regarding the color of certain objects or elements.
@@
  • Text: Questions related to text or symbols present in the image.
@@
SugarCrepe contains three types of negative samples—Add, Replace, and Swap—designed to probe whether vision-language models (VLMs) can recognize text that differs only in subtle structural changes yet conveys a completely different overall meaning in image-text matching tasks.
MLLM benchmarks. We further integrate the fine-tuned vision encoder into LLaVA and evaluate its performance on several MLLM downstream benchmarks, including SEED-Bench (Li et al., 2023a), GQA (Hudson and Manning, 2019), VizWiz (Gurari et al., 2018), PoPE (Li et al., 2023b), TextVQA (Singh et al., 2019), MMBench (Liu et al., 2024b), and VQAv2 (Goyal et al., 2017).
@@
[…]
Table 11: MLLM(7B) performance under $224^{2}$ image resolution.
Method Data Adapter IN-1k I2T Avg. T2I Avg.
CLIP 400M - 74.5 66.9 55.5
[…]
LLM2CLIP 3M $4\timeslinear$ 52.8 82.0 79.1
[…]
LLM2CLIP 15M $4\timeslayers$ 54.0 83.2 80.1
[…]
LLM2CLIP 30M $4\timeslinear$ 56.4 85.8 82.1
[…]
@@
A.3 More results.
Compared to Long-CLIP. Long-CLIP Zhang et al. (2024) is a classic method for extending the long-text capabilities of CLIP. Here, we compare ProCLIP with Long-CLIP. We restricted the data volume for ProCLIP to 1M, consistent with the Long-CLIP experiments. The results are shown in Table 10. Our method outperforms the Long-CLIP paradigm in both short-text and long-text retrieval. Furthermore, we extend CLIP’s capabilities to be both multilingual and fine-grained understanding.
Linear Probe. We conduct linear probe evaluations of the model on 11 datasets. As shown in Tab.7, our method consistently achieves superior performance. This advantage stems from our progressive alignment framework, which stabilizes training through two-stage regularization that prevents overfitting in the vision encoder while preserving generalization capability.
@@
[…]
Table 13: Linear Probe performance on 11 datasets.
@@
A.4 Future Works.
Training Efficiency. ProCLIP is subject to the limitation of increased computational overhead. We consider the following directions to potentially reduce computational overhead:
@@
Fine-grained Visual Alignment. Future work explores the integration of a dedicated local alignment loss to explicitly strengthen the model’s fine-grained visual-textual alignment capabilities, thereby enhancing performance on tasks that require detailed spatial understanding.
More Model Architecture. Our approach replaces the original CLIP text encoder with an LLM-based embedder to enhance multiple capabilities. From another perspective, we consider whether the vision encoder in the dual-tower architecture can also be replaced to address limitations in visual representation. For instance, the CLIP image encoder is known to lack locality; substituting the image encoder may mitigate this limitation. We leave a thorough investigation of this direction for future work.
A.5 Visualization.
As illustrated in Fig. 9, ProCLIP effectively extends long-text processing capabilities, enabling precise image-text alignment even within highly detailed descriptions.
Figure 9: The visualization of ProCLIP vs CLIP in long-text cross-modal retrieval. From top to bottom: Case 1, Case 2, Case 3, and Case 4.